# Linear orders on chainable continua

### v1

@@ Abstract.
We prove that there exist exactly $2$ distinct ultrafilter orders on any arc, exactly $4$ distinct ultrafilter orders on the Warsaw sine curve, and exactly $2^{\mathfrak{c}}$ distinct ultrafilter orders on the Knaster continuum. We study the order type of various chainable continua equipped with an ultrafilter order and prove that a chainable continuum $X$ is Suslinian if and only if for every ultrafilter order $\leq_{\mathcal{U}}^{\mathcal{D}}$ on $X$ the space $X$ with an order topology, generated by the order $\leq_{\mathcal{U}}^{\mathcal{D}}$, is ccc.
We study also descriptive complexity of ultrafilter orders on chainable continua. We prove that the existence of closed ultrafilter order characterizes the arc and we show that for Suslinian chainable continua, any ultrafilter order is both of type $F_{\sigma}$ and $G_{\delta}$. On the other hand, we prove that there is no analytic and no co-analytic ultrafilter order on the Knaster continuum.
@@ 1.1. The ultrafilter orders on chainable continua
Let us mention that the idea of considering ultrafilter orders on chainable continua is due to Jakub Różycki. According to our knowledge, the study of such orders on chainable continua is a new concept, which hasn’t been studied before.
@@ 1.2. Structure of the paper
Finally, in Section 7 we prove a new characterization of an arc - it is the only chainable continuum on which there exists a closed ultrafilter order. We show also that for Suslinean chainable continua any ultrafilter order is both of type $F_{\sigma}$ and $G_{\delta}$. Then we present a proof of a theorem that there is no analytic and no co-analytic (so in particular – no Borel) ultrafilter order on the Knaster continuum.
@@ 2. Background on chainable continua
Recall that a chainable continuum is a compact, connected and metrizable topological space $X$, which satisfies the following property: For a fixed metric $d$, generating topology of $X$, and for every $\varepsilon>0$, $X$ can be covered by a finite chain of open sets $d_{1},d_{2},...,d_{n}$ such that diameter of $d_{i}$ is smaller that $\varepsilon$ for each $i$ (we say that sequence of sets $d_{1},d_{2},...,d_{n}$ is a chain if for each $i,j\in\{1,...,n\}$ we have $d_{i}\cap d_{j}\neq\varnothing\iff|i-j|\leq 1$). The elements of the chain, i.e., the sets $\{d_{i}\}_{i=1}^{n}$, will be called links. Equivalently, chainable continuum is an inverse limit of a sequence of arcs.
We will adopt the convention, used, for example, in Bing’s work [Bi1], that we will denote chains with uppercase letters (e.g., $D,E,F$...), and links of chains with lowercase letters (e.g., $\{d_{i}\}_{i},\{e_{j}\}_{j},\{f_{k}\}_{k}$...).
If $\{D_{n}\}_{n=1}^{\infty}$ is a sequence of chains in the space $X$, then we will use the symbol $d_{i,j}$ to denote the ithi-th link in the jthj-th chain of the sequence $\{D_{n}\}_{n=1}^{\infty}$, i.e. the ithi-th link of the chain $D_{j}=\{d_{1,j},...,d_{i-1,j},d_{i,j},d_{i+1,j},...,d_{k_{j},j}\}$.
@@
If $E=\{e_{1},...,e_{n}\}$ is a chain in continuum $X$ and $mesh(E)<\varepsilon$ for a given $\varepsilon>0$, then we say that $E$ is an $\varepsilon$-chain. Thus, a metric continuum $X$ is chainable if and only if for every $\varepsilon>0$ there exists $\varepsilon$-chain covering $X$.
The following fact will be useful when working with chainable continua.
@@
    There is an infinite sequence of chains $D_{1},D_{2},D_{3},$... such that for every $n\in\mathbb{N}\setminus\{0\}$, the chain $D_{n}$ covers $X$ and $mesh(D_{n})\xrightarrow[]{n\to\infty}0$.
@@
It is worth noting, however, that the class of chain continua is much richer – it can be shown that there exists $\mathfrak{c}$ pairwise non-homeomorphic chainable continua [].
@@ 3. Definitions of ultrafilter orders and some of their basic properties
Note that the order $\leq_{\mathcal{U}}^{\mathcal{D}}$ is an ultraproduct (with respect to the ultrafilter $\mathcal{U}$) of the family of orders $\leq_{D_{n}},n\in\mathbb{N}$.
@@
Below we check that relation $\leq_{\mathcal{U}}^{\mathcal{D}}$ is a linear order on $X$.
@@
  • Transitivity: Suppose that $x\leq_{\mathcal{U}}^{\mathcal{D}}y$ and $y\leq_{\mathcal{U}}^{\mathcal{D}}z$. Then $\{n\in\mathbb{N}:x\leq_{D_{n}}y\}\in\mathcal{U}$ oraz $\{n\in\mathbb{N}:y\leq_{D_{n}}z\}\in\mathcal{U}$, hence $\{n\in\mathbb{N}:x\leq_{D_{n}}y\}\cap\{n\in\mathbb{N}:y\leq_{D_{n}}z\}\in\mathcal{U}$. Since $\{n\in\mathbb{N}:x\leq_{D_{n}}z\}\supseteq\{n\in\mathbb{N}:x\leq_{D_{n}}y\}\cap\{n\in\mathbb{N}:y\leq_{D_{n}}z\}\in\mathcal{U}$, we have $x\leq_{\mathcal{U}}^{\mathcal{D}}z$.
@@
    Suppose that both inequalities $i\leq j$ and $l\leq m$ are strict, i.e. $i<j$ and $l<m$. From condition 3.2 we know that j1j-1 is the smallest possible value of $l$ and $i+1$ is the biggest possible value of $m$. Hence from condition $l<m$ it follows that j1<i+1j-1<i+1, so $j<i+2$. Since we know that $i<j$ it must happen that $j=i+1$. It means that the points $x$ and $y$ are in adjacent links of the chain $D_{n}$. But this is a contradiction with $n\in A$. Obtained contradiction proves that $i=j$ or $l=m$. From condition 3.1 we know that for all $n\in A$ points $x$ and $y$ are in the same link of the $n$-th chain. Hence $\{n\in A:\exists_{~r\leq k_{n}}~x,y\in d_{r,n}\}\in\mathcal{U}$.
    Since $x\neq y$, we know that the set $\{n\in A:\exists_{~r\leq k_{n}}~x,y\in d_{r,n}\}$ is a finite set, so it is not in $\mathcal{U}$ and we obtain a contradiction. It means that under assumption $x\leq_{\mathcal{U}}^{\mathcal{D}}y$ and $y\leq_{\mathcal{U}}^{\mathcal{D}}x$ we obtain $x=y$.
Note that the above definition depends on a fixed sequence of chains $\mathcal{D}=\{D_{n}\}_{n\in\mathbb{N}}$ and on a fixed non-principal ultrafilter $\mathcal{U}$ on $\mathbb{N}$. This means that different choices of a sequence of chains covering $X$ or a non-principal ultrafilter on $\mathbb{N}$ can generate different orders on the continuum $X$.
@@
For any non-principal ultrafilter on $\mathbb{N}$ and for any sequence of chains $\mathcal{D}=\{D_{n}\}_{n\in\mathbb{N}}$ covering the chainable continuum $X$, such that $mesh(D_{n})\xrightarrow[]{n\to\infty}0$, the order $\leq_{\mathcal{U}}^{\mathcal{D}}$ is a linear order on $X$.
@@
We will use easy-to-prove lemma from book [Ma].
@@
Let $E_{n}=\{e_{1,n},...,e_{k_{n},n}\}$ be a chain of intervals of diameter at most $\delta_{n}$, numbered according to the standard order on $[0,1]$, covering $I_{n}$. Let $\mathcal{A}_{n}=\{p_{n}^{-1}(e_{i,n})\}_{i=1}^{k_{n}}$. Since $\lim_{n\to\infty}\varepsilon_{n}=0$, then $mesh(\mathcal{A}_{n})\xrightarrow[]{n\to\infty}0$. For $i\leq k_{n}$, let $d_{i,n}=h\circ p_{n}^{-1}(e_{i,n})$ and let $D_{n}=\{d_{i,n}:i\leq k_{n}\}$. We know that the homeomorphism $h$ is uniformly continuous, so
@@
Thus we have proved that any ultrafilter order on $Y$ generates an ultrafilter order on $X$. However we don’t know if the converse holds, i.e. if it is true that for a chainable continuum $X$ and for the inverse limit of a sequence of arcs $Y$, homeomorphic to $X$, any ultrafilter order on $X$ generates some ultrafilter order on $Y$ (see Question 8.1).
@@ 4.1. Ultrafilter orders on an arc
Let $X$ be a chainable continuum and let $x,y,z\in X$. Let $\mathcal{D}=\{D_{n}\}_{n\in\mathbb{N}}$ be any sequence of chains covering $X$, such that $mesh(D_{n})\xrightarrow[]{n\to\infty}0$. Suppose there exists a continuum $M\subseteq X$ such that $x,y\in M$ and $z\notin M$. Then
@@
Let us assume that $\{T_{i}:i\in I\}$ are the arc components of a chainable continuum $X$ and that for some $i\in I,$ $x,y\in T_{i}$. Suppose that $j\neq i$ and $z\in T_{j}$. Since $x$ and $y$ are in the same arc component of $X$, we can connect them with an arc that does not contain $z$ and then use Theorem 4.2. Thus the following holds:
@@ 4.2. Ultrafilter orders on continua $S_{1},S_{2}$ and $S_{3}$
Now we prove that we can define at least four distinct ultrafilter orders on $S_{1}$. There exist sequences of chains
@@
    for for each $n\in\mathbb{N}$, the point (0,1)(0,-1) belongs to the $m^{\prime}$-th link $d^{\prime}_{m^{\prime},n}$ of the chain $D^{\prime}_{n}=\{d^{\prime}_{i,n}\}_{i=1}^{k^{\prime}_{n}}$, where m’=min{i ≤k’_n: d’_i,n∩({0} ×[-1,1]) ≠∅},
@@
    for each $n\in\mathbb{N}$, the chain $E_{n}\in\mathcal{E}$ is the reversely numbered chain $D_{n}$ (i.e., E_n = {e_i,n}_i=1^k_n and for i ∈{1,…,k_n}  e_i,n = d_k_n-i+1, n ),
@@
    for each $n\in\mathbb{N}$ the chain $E^{\prime}_{n}\in\mathcal{E}^{\prime}$ is the reversely numbered chain $D^{\prime}_{n}$ (i.e., E’_n = {e’_i,n}_i=1^k’_n and for i ∈{1,…,k’_n}  e’_i,n = d’_k’_n-i+1, n ).
@@
$(0,1)<_{\mathcal{U}}^{\mathcal{D}}(0,-1),~\Big(\frac{2}{7\pi},-1\Big)>_{\mathcal{U}}^{\mathcal{D}}\Big(\frac{2}{3\pi},-1\Big),~(0,1)>_{\mathcal{U}}^{\mathcal{D}^{\prime}}(0,-1),~\Big(\frac{2}{7\pi},-1\Big)>_{\mathcal{U}}^{\mathcal{D}^{\prime}}\Big(\frac{2}{3\pi},-1\Big)$
$(0,1)>_{\mathcal{U}}^{\mathcal{E}}(0,-1),~\Big(\frac{2}{7\pi},-1\Big)<_{\mathcal{U}}^{\mathcal{E}}\Big(\frac{2}{3\pi},-1\Big),~(0,1)<_{\mathcal{U}}^{\mathcal{E}^{\prime}}(0,-1),~\Big(\frac{2}{7\pi},-1\Big)<_{\mathcal{U}}^{\mathcal{E}^{\prime}}\Big(\frac{2}{3\pi},-1\Big)$
@@
$S_{3}=\bigcup_{n\in\mathbb{N}\setminus\{0\}}I_{n}\cup\bigcup_{n\in\mathbb{N}\setminus\{0\}}A_{n}\cup\{(0,0)\},$
@@
Let $n,m\in\mathbb{N}$ and let $x\in\{0,1\}^{\mathbb{N}}$. Denote $I_{n}=\{\frac{1}{n}\}\times[0,1]$ and let $D_{m}=\{d_{1,m},...,d_{k_{m},m}\}$ be a chain covering the continuum $S_{3}$. We say that the chain $D_{m}$ covers the interval $I_{n}$ according to the $n$-th coordinate $x_{n}$ of the sequence $x$ if the following condition holds:
@@
The following illustrations (Figure 3, Figure 4 and Figure 5) show an example of selecting the chains $D_{1}=\{d_{i}\}_{i=1}^{10},D_{2}=\{d_{i}\}_{i=1}^{13},D_{3}=\{d_{i}\}_{i=1}^{19}$ for the sequence x=(0,$1,$)x=(0,$1,$...) such that
@@
    for each $n\in\mathbb{N}\setminus\{0\}$, $mesh(D_{n})<\frac{1}{n},$
@@
    for each $n\in\mathbb{N}\setminus\{0\}$, for each $i\leq n$ the chain $D_{n}=\{d_{1,n},...,d_{k_{n},n}\}$ covers the interval $I_{i}=\{\frac{1}{i}\}\times[0,1]$ according to the $i$-th coordinate $x_{i}$ of the sequence $x$.
Let $x,y\in\{0,1\}^{\mathbb{N}}$ be distinct binary sequences, and $\mathcal{D}_{x}=\{D_{x,n}\}_{n\in\mathbb{N}},\\ \mathcal{D}_{y}=\{D_{y,n}\}_{n\in\mathbb{N}}$ be sequences of chains assigned to the sequences $x$ and $y$, respectively, satisfying two above conditions. Let $n\in\mathbb{N}$ be such that $x_{n}\neq y_{n}$. Without loss of generality, assume that $x_{n}=0$ and $y_{n}=1$. From condition 2, we know that for $k\geq n$, the chain $D_{x,k}\in\mathcal{D}_{x}$ covers $I_{n}$ according to $x_{n}$, and the chain $D_{y,k}\in\mathcal{D}_{y}$ covers $I_{n}$ according to $y_{n}$. Therefore, for $k\geq n$, we have
@@
Therefore, the orders $\leq_{\mathcal{U}}^{\mathcal{D}_{x}}$ and $\leq_{\mathcal{U}}^{\mathcal{D}_{y}}$ are different. We know that $|\{0,1\}^{\mathbb{N}}|=\mathfrak{c}$. Thus, we have proved that there exists $\mathfrak{c}$ pairwise distinct ultrafilter orders on $S_{3}$.
@@
Let $\{T_{i}:i\in I\}$ be the arc components of a chainable continuum $X$. Notice that by Corollary 4.4 we know that for all $i\neq j$ we have (see Definition 3.1)

### v2

@@ Abstract.
We prove that there exist exactly $2$ distinct ultrafilter orders on any arc, exactly $4$ distinct ultrafilter orders on the Warsaw sine curve, and exactly $2^{\mathfrak{c}}$ distinct ultrafilter orders on the Knaster continuum. We study the order type of various chainable continua equipped with an ultrafilter order and prove that a chainable continuum $X$ is Suslinean if and only if for every ultrafilter order $\leq_{\mathcal{U}}^{\mathcal{D}}$ on $X$, the space $(X,\leq_{\mathcal{U}}^{\mathcal{D}})$ is order isomorphic to $([0,1],\leq)$.
We study also descriptive complexity of ultrafilter orders on chainable continua. We prove that the existence of closed ultrafilter order characterizes the arc and we show that for Suslinean chainable continua, any ultrafilter order is both of type $F_{\sigma}$ and $G_{\delta}$. On the other hand, we prove that there is no analytic and no co-analytic ultrafilter order on the Knaster continuum.
@@ 1. Introduction
Convention
Throughout the text, we will use the convention that the set of natural numbers is equal to the set of positive integers, i.e. $\mathbb{N}=\{1,2,3,$...\}.
@@ 1.1. The ultrafilter orders on chainable continua
Let us mention that the idea of considering ultrafilter orders on chainable continua is due to Jakub Różycki. According to our knowledge, the study of such orders on chainable continua is a new concept, which has not been studied before.
@@ 1.2. Structure of the paper
In Section 7 we prove a new characterization of an arc - it is the only chainable continuum on which there exists a closed ultrafilter order. We show also that for Suslinean chainable continua any ultrafilter order is both of type $F_{\sigma}$ and $G_{\delta}$. Then we present a proof of a theorem that there is no analytic and no co-analytic (so in particular – no Borel) ultrafilter order on the Knaster continuum.
Finally, in Section 8 we study the relationships between the endpoints and the absolute endpoints of chainable continua and the minimal or maximal points of chainable continua equipped with ultrafilter orders.
@@ 2. Background on chainable continua
Recall that a chainable continuum is a compact, connected and metrizable topological space $X$, which satisfies the following property: For a fixed metric $d$, generating topology of $X$, and for every $\varepsilon>0$, $X$ can be covered by a finite chain of open sets $d_{1},d_{2},...,d_{n}$ such that diameter of $d_{i}$ is smaller that $\varepsilon$ for each $i$ (we say that sequence of sets $d_{1},d_{2},...,d_{n}$ is a chain if for each $i,j\in\{1,...,n\}$ we have $d_{i}\cap d_{j}\neq\varnothing\iff|i-j|\leq 1$). The elements of the chain, i.e., the sets $\{d_{i}\}_{i=1}^{n}$, will be called links. Equivalently, a chainable continuum is an inverse limit of a sequence of arcs.
We will use the convention from Bing’s work [Bi1] – we will denote chains with uppercase letters (e.g., $D,E,F$...), and links of chains with lowercase letters (e.g., $\{d_{i}\}_{i},\{e_{j}\}_{j},\{f_{k}\}_{k}$...).
If $\{D_{n}\}_{n=1}^{\infty}$ is a sequence of chains in the space $X$, then we will use the symbol $d_{i,j}$ to denote the ithi-th link in the jthj-th chain of the sequence $\{D_{n}\}_{n=1}^{\infty}$, i.e. the $i$-th link of the chain $D_{j}=\{d_{1,j},...,d_{i-1,j},d_{i,j},d_{i+1,j},...,d_{k_{j},j}\}$.
@@
If $E=\{e_{1},...,e_{n}\}$ is a chain in continuum $X$ and $mesh(E)<\varepsilon$ for a given $\varepsilon>0$, then we say that $E$ is an $\varepsilon$-chain. Thus, a metric continuum $X$ is chainable if and only if for every $\varepsilon>0$ there exists an $\varepsilon$-chain covering $X$.
The following fact will be useful to work with chainable continua.
@@
    There is an infinite sequence of chains $D_{1},D_{2},D_{3},$... such that for every $n\in\mathbb{N}$, the chain $D_{n}$ covers $X$ and $mesh(D_{n})\xrightarrow[]{n\to\infty}0$.
@@
It is worth noting, however, that the class of chainable continua is much richer – it can be shown that there exists $\mathfrak{c}$ pairwise non-homeomorphic chainable continua [].
@@ 3. Definitions of ultrafilter orders and some of their basic properties
Note that the order $\leq_{\mathcal{U}}^{\mathcal{D}}$ is an ultraproduct (with respect to the ultrafilter $\mathcal{U}$) of the family of relations $\leq_{D_{n}},n\in\mathbb{N}$.
@@
Below we check that relation $\leq_{\mathcal{U}}^{\mathcal{D}}$ is an order on $X$.
@@
  • Transitivity: Suppose that $x\leq_{\mathcal{U}}^{\mathcal{D}}y$ and $y\leq_{\mathcal{U}}^{\mathcal{D}}z$. Then $\{n\in\mathbb{N}:x\leq_{D_{n}}y\}\in\mathcal{U}$ and $\{n\in\mathbb{N}:y\leq_{D_{n}}z\}\in\mathcal{U}$, hence $\{n\in\mathbb{N}:x\leq_{D_{n}}y\}\cap\{n\in\mathbb{N}:y\leq_{D_{n}}z\}\in\mathcal{U}$. Since $\{n\in\mathbb{N}:x\leq_{D_{n}}z\}\supseteq\{n\in\mathbb{N}:x\leq_{D_{n}}y\}\cap\{n\in\mathbb{N}:y\leq_{D_{n}}z\}\in\mathcal{U}$, we have $x\leq_{\mathcal{U}}^{\mathcal{D}}z$.
@@
    Suppose that both inequalities $i\leq j$ and $l\leq m$ are strict, i.e. $i<j$ and $l<m$. From condition 3.2 we know that j1j-1 is the smallest possible value of $l$ and $i+1$ is the biggest possible value of $m$. Hence from condition $l<m$ it follows that j1<i+1j-1<i+1, so $j<i+2$. Since we know that $i<j$ it must happen that $j=i+1$. It means that the points $x$ and $y$ are in adjacent links of the chain $D_{n}$. But this is a contradiction with $n\in A$. This proves that $i=j$ or $l=m$. From condition 3.1 we know that for all $n\in A$ points $x$ and $y$ are in the same link of the $n$-th chain. Hence $\{n\in A:\exists_{~r\leq k_{n}}~x,y\in d_{r,n}\}\in\mathcal{U}$.
    Since $x\neq y$, we know that the set $\{n\in A:\exists_{~r\leq k_{n}}~x,y\in d_{r,n}\}$ is finite, so it is not in $\mathcal{U}$ and we obtain a contradiction. It means that under assumption $x\leq_{\mathcal{U}}^{\mathcal{D}}y$ and $y\leq_{\mathcal{U}}^{\mathcal{D}}x$ we obtain $x=y$.
Note that Definition 3.2 depends on a fixed sequence of chains $\mathcal{D}=\{D_{n}\}_{n\in\mathbb{N}}$ and on a fixed non-principal ultrafilter $\mathcal{U}$ on $\mathbb{N}$. This means that different choices of a sequence of chains covering $X$ or a non-principal ultrafilter on $\mathbb{N}$ can generate different orders on the continuum $X$.
@@
For any non-principal ultrafilter on $\mathbb{N}$ and for any sequence of chains $\mathcal{D}=\{D_{n}\}_{n\in\mathbb{N}}$ covering a chainable continuum $X$, such that $mesh(D_{n})\xrightarrow[]{n\to\infty}0$, the order $\leq_{\mathcal{U}}^{\mathcal{D}}$ is a linear order on $X$.
@@
We will use an easy-to-prove lemma from book [Ma].
@@
Let $E_{n}=\{e_{1,n},...,e_{k_{n},n}\}$ be a chain of intervals of diameter at most $\delta_{n}$, enumerated according to the standard order on $[0,1]$, covering $I_{n}$. Let $\mathcal{A}_{n}=\{p_{n}^{-1}(e_{i,n})\}_{i=1}^{k_{n}}$. Since $\lim_{n\to\infty}\varepsilon_{n}=0$, then $mesh(\mathcal{A}_{n})\xrightarrow[]{n\to\infty}0$. For $i\leq k_{n}$, let $d_{i,n}=h\circ p_{n}^{-1}(e_{i,n})$ and let $D_{n}=\{d_{i,n}:i\leq k_{n}\}$. We know that the homeomorphism $h$ is uniformly continuous, so
@@
Thus we have proved that any ultrafilter order on $Y$ generates an ultrafilter order on $X$. However we don’t know if the converse holds, i.e. if it is true that for a chainable continuum $X$ and for the inverse limit of a sequence of arcs $Y$, homeomorphic to $X$, any ultrafilter order on $X$ generates some ultrafilter order on $Y$ (see Question 9.1).
@@ 4.1. Ultrafilter orders on an arc
Let $X$ be a chainable continuum and let $x,y,z\in X$. Let $\mathcal{D}=\{D_{n}\}_{n\in\mathbb{N}}$ be any sequence of chains covering $X$, such that $mesh(D_{n})\xrightarrow[]{n\to\infty}0$. Suppose that there exists a continuum $M\subseteq X$ such that $x,y\in M$ and $z\notin M$. Then
@@
Let us assume that $\{T_{i}:i\in I\}$ are the arc components of a chainable continuum $X$ and that for some $i\in I,$ $x,y\in T_{i}$. Suppose that $j\neq i$ and $z\in T_{j}$. Since $x$ and $y$ are in the same arc component of $X$, we can connect them with an arc that does not contain $z$ and then use Lemma 4.2. Thus the following holds:
@@ 4.2. Ultrafilter orders on continua $S_{1},S_{2}$ and $S_{3}$
Now we prove that we can find at least four distinct ultrafilter orders on $S_{1}$. There exist sequences of chains
@@
    for each $n\in\mathbb{N}$, the point (0,1)(0,-1) belongs to the $m^{\prime}$-th link $d^{\prime}_{m^{\prime},n}$ of the chain $D^{\prime}_{n}=\{d^{\prime}_{i,n}\}_{i=1}^{k^{\prime}_{n}}$, where m’=min{i ≤k’_n: d’_i,n∩({0} ×[-1,1]) ≠∅},
@@
    for each $n\in\mathbb{N}$, the chain $E_{n}\in\mathcal{E}$ is the reversely enumerated chain $D_{n}$ (i.e., E_n = {e_i,n}_i=1^k_n and for i ∈{1,…,k_n}  e_i,n = d_k_n-i+1, n ),
@@
    for each $n\in\mathbb{N}$ the chain $E^{\prime}_{n}\in\mathcal{E}^{\prime}$ is the reversely enumerated chain $D^{\prime}_{n}$ (i.e., E’_n = {e’_i,n}_i=1^k’_n and for i ∈{1,…,k’_n}  e’_i,n = d’_k’_n-i+1, n ).
@@
$(0,1)<_{\mathcal{U}}^{\mathcal{D}}(0,-1),~\Big(\frac{2}{7\pi},-1\Big)>_{\mathcal{U}}^{\mathcal{D}}\Big(\frac{2}{3\pi},-1\Big),~(0,1)>_{\mathcal{U}}^{\mathcal{D}^{\prime}}(0,-1),~\Big(\frac{2}{7\pi},-1\Big)>_{\mathcal{U}}^{\mathcal{D}^{\prime}}\Big(\frac{2}{3\pi},-1\Big),$
$(0,1)>_{\mathcal{U}}^{\mathcal{E}}(0,-1),~\Big(\frac{2}{7\pi},-1\Big)<_{\mathcal{U}}^{\mathcal{E}}\Big(\frac{2}{3\pi},-1\Big),~(0,1)<_{\mathcal{U}}^{\mathcal{E}^{\prime}}(0,-1),~\Big(\frac{2}{7\pi},-1\Big)<_{\mathcal{U}}^{\mathcal{E}^{\prime}}\Big(\frac{2}{3\pi},-1\Big).$
@@
$S_{3}=\bigcup_{n\in\mathbb{N}}I_{n}\cup\bigcup_{n\in\mathbb{N}}A_{n}\cup\{(0,0)\},$
@@
Let $n,m\in\mathbb{N}$ and let $x\in\{0,1\}^{\mathbb{N}}$. Denote $I_{n}=\{\frac{1}{n}\}\times[0,1]$ and let $D_{m}=\{d_{1,m},...,d_{k_{m},m}\}$ be a chain covering continuum $S_{3}$. We say that the chain $D_{m}$ covers the interval $I_{n}$ according to the $n$-th coordinate $x_{n}$ of the sequence $x$ if the following condition holds:
@@
The following illustrations (Figure 3, Figure 4 and Figure 5) show an example of a choice of the chains $D_{1}=\{d_{i}\}_{i=1}^{10},D_{2}=\{d_{i}\}_{i=1}^{13},D_{3}=\{d_{i}\}_{i=1}^{19}$ for the sequence x=(0,$1,$)x=(0,$1,$...) such that
@@
    for each $n\in\mathbb{N}$, $mesh(D_{n})<\frac{1}{n},$
@@
    for each $n\in\mathbb{N}$, for each $i\leq n$ the chain $D_{n}=\{d_{1,n},...,d_{k_{n},n}\}$ covers the interval $I_{i}=\{\frac{1}{i}\}\times[0,1]$ according to the $i$-th coordinate $x_{i}$ of the sequence $x$.
Let $x,y\in\{0,1\}^{\mathbb{N}}$ be a distinct binary sequences, and let $\mathcal{D}_{x}=\{D_{x,n}\}_{n\in\mathbb{N}},\\ \mathcal{D}_{y}=\{D_{y,n}\}_{n\in\mathbb{N}}$ be sequences of chains assigned to the sequences $x$ and $y$, respectively, satisfying two above conditions. Let $n\in\mathbb{N}$ be such that $x_{n}\neq y_{n}$. Without loss of generality, assume that $x_{n}=0$ and $y_{n}=1$. From condition 2, we know that for $k\geq n$, the chain $D_{x,k}\in\mathcal{D}_{x}$ covers $I_{n}$ according to $x_{n}$, and the chain $D_{y,k}\in\mathcal{D}_{y}$ covers $I_{n}$ according to $y_{n}$. Therefore, for $k\geq n$, we have
@@
Therefore, the orders $\leq_{\mathcal{U}}^{\mathcal{D}_{x}}$ and $\leq_{\mathcal{U}}^{\mathcal{D}_{y}}$ are distinct.
@@
Let $\{T_{i}\}_{i=1}^{3}$ be the arc components of a chainable continuum $X$. Notice that by Corollary 4.4 we know that for all $i\neq j$ we have (see Definition 3.1)